# literature.bib

### v1

@@ 5 Benchmarking
Throughout Section 3, we have changed several aspects of the AD implementation in order to enable hybrid parallel discrete adjoints. To understand the impact of these changes on the overall hybrid AD performance as well as the impact of the optimizations described in Section 4, we conduct detailed studies of the single-socket performance on two test cases in Section 5.1. In Section 5.2, we study the impact of the loop parallelization strategies discussed in Section 4. Finally, we assess the scalability with a large test case in Section 5.3. In all test cases, we start with converged primal solutions of the RANS equations with the Spalart-Allmaras (SA) turbulence model [SpalartA1992]. The fluid is modelled as an ideal gas. We discretize convective fluxes with the JST scheme [JamesonST1981] and compute spatial gradients with the Green-Gauss approach. See [Blazek2015] for an overview over the governing equations and numerical schemes. We configure discrete adjoints to compute derivatives of the aerodynamic lift with respect to mesh coordinates, which could, for example, be used for shape optimization. Our goal is to assess comparable computational workloads across different parallel setups, and we therefore always execute a fixed number of discrete adjoint iterations. In particular, we do not make statements about discrete adjoint convergence. Unless mentioned otherwise, we parallelize loops with edge coloring, using the default edge color group size of 512 and adaptivity for discrete adjoints.
@@
Figure 5.1: Decomposition of the discrete adjoint wall clock time.
We decompose the wall clock time of discrete adjoint computations as visualized in Fig. 5.1. We distinguish between AD-relevant timings on the one hand, that is, all timings that are directly influenced by the changes described in Sections 3 and 4, and non-AD timings on the other hand, like preprocessing of the geometry or file I/O. While non-AD timings might be different between MPI-only and hybrid builds of SU2, this is not influenced by hybrid AD itself and not of primary interest in our comparisons. We further refine the AD-relevant timings into recording, evaluation, and management. Recording encompasses the primary and secondary recording passes as explained in Section 2 as well as passive computations for the purpose of clearing identifiers. Evaluation refers to all tape evaluations in the course of discrete adjoint iterations (2.2) or (2.4), respectively, as well as the final evaluation with respect to parameters in (2.3). Management encompasses interactions with CoDiPack’s vector of adjoint variables, that is, the seeding prior to tape evaluations as well as the extraction after tape evaluations.
Support for hybrid parallel discrete adjoints is available in the upstream version of SU2. We make a few changes to ensure comparable computational workloads across different parallel setups (e. g., no special behaviour for the case of a single OpenMP thread, lower tolerances so that iterations do not stop early). We use release builds of SU2, compiled with GCC 11.3, with MPI support (OpenMPI 4.0), with AD enabled, with AVX512, and use single precision for linear algebra. The changes made to SU2 and precise build settings are documented in the hybrid_parallel_discrete_adjoints branch [^6] https://github.com/jblueh/SU2/tree/hybrid_parallel_discrete_adjoints, alongside the test cases that are used in the following [^7] https://github.com/jblueh/SU2/tree/hybrid_parallel_discrete_adjoints/benchmark.
@@ 5.1 Single-Socket Performance
Figure 5.2: NACA 0012 mesh (left) and Onera M6 mesh (right) for single-socket performance studies.
@@
    We consider the flow around a 2D NACA 0012 airfoil at Mach 0.5 and $0^{\circ}$ angle of attack (AoA), with a Reynolds number (Re) of $20\cdot 10^{6}$. We assume a constant dynamic viscosity of $1.02\cdot 10^{-5}\,\text{Pa\,s}$. The computational domain is discretized by means of a structured mesh with 75140 quadrilaterals, as displayed in the left part of Fig. 5.2. Given the converged primal flow solution, we perform 300 discrete adjoint iterations in the spirit of (2.2) followed by one evaluation of (2.3).
@@
To understand the differences in the AD-specific timings in detail, we consider various intermediate build configurations of SU2 that illustrate the key steps in the transition to hybrid parallel discrete adjoints. We begin with the classical, MPI-only build of SU2 (MPI build, linear). We first exchange the linear identifier management scheme with a reuse scheme (MPI build, reuse). Next, we disable preaccumulations that are incompatible with hybrid parallel AD (MPI build, reuse, hybr. preacc.). We then transition to a hybrid build that uses OpDiLib and features all changes described in Section 3 (hybrid build, no opt.). Finally, we enable the optimizations discussed in Section 4, first only optimized adjoint vector management (hybrid build, opt. adj. mgmt.) and then in addition also the shared reading optimization (hybrid build, both opt.).
Figure 5.3: AD-specific performance of the NACA 0012 test case (top) and the Onera M6 test case (bottom). Recording performance (left), management performance (middle), and evaluation performance (right). Serial and parallel timings for various build configurations. Error bars indicate the variation across runs. Speedup factors are relative to the serial run of the respective build.
The left part of Fig. 5.3 displays the AD recording performance. The respective serial recording timings, for both the NACA 0012 and the Onera M6 test case, illustrate the key differences between the builds. Switching to a reuse management scheme incurs performance overhead due to increased tape size and more involved identifier management, as explained in detail in [SagebaumBG2020]. Enabling only preaccumulations compatible with hybrid AD reduces the overall number of preaccumulations, which usually reduces the recording time slightly. Switching to a hybrid build requires a thread-safe CoDiPack type, which incurs costs mainly due to tapes accessed via static thread-local pointers [BluehdornSG2023]. Optimizing for adjoint vector management makes the preaccumulations throughout the recording more efficient, also in the serial case. The shared reading optimization has only a minor impact on the recording performance, the observed small overheads could be due to the tracking of code parts suitable for evaluation without atomics.
@@
The right part of Fig. 5.3 showcases the AD evaluation performance of the two test cases, which both show similar trends. We explain the evolution of the serial performance. Reuse management leads to larger tapes, both in terms of tape entry size and the overall number of tape entries, and correspondingly larger evaluation times [SagebaumBG2020]. Likewise, disabling some of the preaccumulations shifts workload from the recording phase to the evaluation phase. The hybrid build’s serial evaluation performance demonstrates the absolute cost of performing atomic updates on adjoint variables. While optimized adjoint vector management does not affect the evaluation performance, the shared reading optimization reduces the overhead by approximately fifty percent in both cases, which underlines the importance of eliminating atomic updates on adjoint variables. We observe consistent MPI speedups of ×$\times 10.3$ to ×$\times 10.9$ and ×$\times 9.3$ to ×$\times 10.4$ for the NACA 0012 and the Onera M6 test case, respectively, and respective OpenMP speedups in the ranges ×$\times 9.5$ to ×$\times 9.9$ and ×$\times 9.8$ to ×$\times 10.0$. As the evaluation inherits the parallelism of the recording, serial parts that limit the OpenMP scaling of the recording will limit the scaling of the evaluation as well. In addition, serialization due to atomic updates on adjoint variables could be a limiting factor for the OpenMP speedup. The observed OpenMP speedups are competitive, which is important for the overall performance as the evaluation time scales with the number of discrete adjoint iterations.
@@
Figure 5.4: NACA 0012 performance (top) and Onera M6 performance (bottom). Memory high-water marks depending on the type and degree of parallelism for the five SU2 build configurations (left), and joint memory consumption and evaluation time for selected configurations with varying degrees of parallelism (right).
@@
We investigate the memory consumption of the build configurations with varying degrees of parallelism for both test cases in the left part of Fig. 5.4. We first explain the differences between the build configurations at serial execution. When we compare the memory consumption of hybrid builds to the memory consumption of the classical MPI-only build with linear management, we see that one part of the difference is due to the switch to reuse management, which requires larger individual tape entries and, as copy operations need to be recorded, at the same time more overall tape entries, see [SagebaumBG2020], where the benefits of optimized copies for SU2 were observed as well. Another part is due to disabling incompatible preaccumulations, which results in larger tapes and therefore higher memory consumption. While both types of overhead have similar magnitudes for the NACA 0012 test case, the additional memory consumption due to disabled preaccumulations becomes dominant in the Onera M6 test case. The optimizations from Section 4 do not affect the memory consumption. For each test case, the memory usage increases linearly with an increasing number of MPI processes with similar trends across all build configurations. We attribute this to data duplication across processes, additional halo layers in the partitioning, and larger communication buffers. The memory overhead of increasing the number of OpenMP threads, on the other hand, is significantly smaller, so much that for the NACA 0012 test case, OpenMP parallel execution of hybrid builds requires less per-node memory than MPI parallel execution of the classical build configuration for this test case already at 10-fold parallelism, despite the larger base memory consumption. For the Onera M6 test case, on the other hand, 12-fold parallelism does not suffice to compensate for the memory overhead due to disabled preaccumulations.
@@
We assess the performance tradeoffs in terms of both memory and AD-specific time (aggregated time spent on recording, management, and evaluation), which we plot for selected configurations and their varying degrees of parallelism in the right part of Fig. 5.4. We consider the classical MPI build with linear identifier management and the fully optimized hybrid build. Clearly, the former dominates the latter in terms of both memory and runtime for the Onera M6 test case. While the hybrid capabilities allow trading memory for runtime at higher degrees of parallelism in the NACA 0012 test case, the MPI build with linear identifier management at 8-fold parallelism still dominates higher degrees of hybrid parallelism. This suggests that, without further improvements, the classical MPI parallelism is the preferred choice for single-socket discrete adjoint computations in SU2. The same recommendation was already made regarding the choice of parallelism for primal single-socket computations in SU2 [GomesEP2021]. In the right part of Fig. 5.4, we also consider the MPI build with reuse identifier management and only hybrid AD compatible preaccumulations, which is the MPI build that is most similar to the hybrid build in terms of AD and admits a fair assessment of the hybrid AD performance. Starting already at small degrees of parallelism, we see that hybrid AD can serve as a tool to balance memory consumption and runtime.
@@
The results also broaden the perspective on the hybrid parallel discrete adjoint performance. For instance, we expect the notable difference in the memory scaling (left part of Fig. 5.4) to continue for larger numbers of cores, which could make hybrid parallel discrete adjoints attractive on CPUs with larger NUMA domains. Any principal improvement of hybrid parallel AD such as reduction of the memory offset or improvements of the evaluation time could bridge the small margin between the classical MPI build and the fully optimized hybrid build for the NACA 0012 test case (top right part of Fig. 5.4). Finally, as the memory overhead in the Onera M6 test case (bottom left part of Fig. 5.4) is mostly due to disabling preaccumulations, we identify them as a promising target for further improvements of the hybrid parallel discrete adjoint performance.
@@ 5.2 Loop Parallelization Strategies
We conduct a performance study to support the design decisions regarding loop parallelization for discrete adjoints as outlined in Section 4. We study the performance of the Onera M6 test case described in Section 5.1 with both the reduction strategy and edge coloring for loop parallelization, where we vary the edge color group size in the latter. We also assess the performance of the adaptive choice of the edge color group size that we motivated in Section 4. The results are displayed in Fig. 5.5.
Figure 5.5: Performance of the Onera M6 test case. We compare reductions (violet) to edge coloring with various edge color group sizes (teal and blue). Our adaptive algorithm chooses an edge color group size of 519 (blue). The performance plots show physical memory high-water marks (top left), evaluation time (top right), and recording time (bottom left) together with the variation across multiple runs. The efficiency measure (bottom right) quantifies the imbalance resulting from edge color groups, with an efficiency threshold of 0.875 (red).
@@ 5.3 Scalability
Figure 5.6: HL-CRM mesh for scaling studies (left) and HL-CRM surface sensitivity (right). The latter is obtained from the derivative of the aerodynamic lift with respect to all mesh coordinates in an additional projection step, treating only surface nodes as design variables.
To assess the scalability of the AD-specific performance, we use an official test case of the Third AIAA CFD High Lift Prediction Workshop involving NASA’s High Lift Common Research Model (HL-CRM) [RumseySS2019], with $8^{\circ}$ AoA at Mach 0.2, $\text{Re}=3.26\cdot 10^{6}$ and a Reynolds length of 275.8 inches. Like in the Onera M6 test case, the dynamic viscosity follows Sutherland’s law with parameters for air. This test case was also used to assess SU2’s hybrid parallel primal performance in [GomesEP2021]. We work with the coarse version of an unstructured mixed-element HL-CRM mesh that consists of approximately 18 million cells and 8.3 million nodes, displayed together with the surface sensitivity in Fig. 5.6. We start with the converged primal solution and solve Eq. 2.4 with 1000 GMRES iterations, restarting the Krylov solver every 100th iteration. Afterwards, we evaluate (2.3) once.
@@
Figure 5.7: AD-specific performance of the HL-CRM test case. Recording performance (left), management performance (middle), and evaluation performance (right). Parallel timings for various build configurations, ranging from 192-fold to 768-fold parallelism. Error bars indicate the variation across runs. Speedups are relative to the respective timings at 192-fold parallelism.
@@
Fig. 5.7 showcases the recording, management, and evaluation performance for the HL-CRM test case for the build configurations with various degrees of parallelism. The trends agree with the previous observations in Section 5.1. Considering the recording performance, the relative difference between MPI builds and the hybrid build stands out in this plot. However, this difference is actually similar or smaller than the corresponding differences in the Onera M6 test case (comparing 12-fold MPI parallelism with 12-fold OpenMP parallelism). Even though the recording is not the major part of the AD-specific runtime, it could be worthwhile to identify and, if possible, improve its parts that do not scale with OpenMP. We see that sufficiently large degrees of parallelism close the gap in the management performance between the hybrid build and the most similar non-hybrid build.
@@
Figure 5.8: HL-CRM memory high-water marks for the four builds and varying degree of parallelism (left), and joint memory consumption and evaluation time for selected configurations with varying degrees of parallelism (right). Plots display joint memory usage across all involved nodes.
@@
The left part of Fig. 5.8 displays the memory consumption of the HL-CRM test case with the different build configurations and depending on the degree of parallelism. The plot clearly shows that reuse identifier management does not only introduce an offset in the memory consumption (as observed in the previous test cases) but also affects the scaling of the memory usage. More MPI processes lead to a larger overall number of halo cells and a larger overall number of copy operations for the associated MPI buffers. Unlike recordings with linear identifier management, approaches with reuse identifier management have to record copy operations on the tape, which could explain why their memory usage grows faster with an increasing number of MPI processes. Even though hybrid builds rely on a reuse management scheme as well, introducing 12-fold OpenMP parallelism per MPI process is still sufficient to achieve better memory scaling than the MPI build with linear identifier management. The crossover point lies at approximately 384-fold parallelism for this test case. If we compare the memory consumption of the hybrid configuration with the most similar non-hybrid configuration (MPI, reuse, hybr. preacc.), we see that the hybrid AD approach can save substantial amounts of memory, there is a reduction by a factor of approximately ×$\times 2.4$ at 768-fold parallelism. The plot also shows the additional memory consumption due to disabling preaccumulations, clearly visible as the offset between the curves for MPI builds with reuse management and enabled and disabled preaccumulations, respectively. Re-enabling preaccumulations in the hybrid build could improve its memory consumption up to the same offset. It would shift the crossover point to notably smaller degrees of parallelism and approximately double the memory savings at 768-fold parallelism due to moving from the MPI build with linear management and MPI parallel execution to a hybrid build with hybrid parallel execution in this test case. We expect further improvements on CPU with larger NUMA domains that admit higher degrees of OpenMP parallelism per MPI process.
@@
The right part of Fig. 5.8 summarizes the performance in terms of both memory consumption and AD-specific runtime. If we compare the hybrid configuration with the most similar non-hybrid configuration in terms of AD, we see that hybrid AD can be used to trade memory for runtime also in this large test case and correspondingly higher degrees of parallelism. While there are similar tradeoffs between the hybrid configuration and the MPI configuration with linear management, the runtime price for memory reductions is relatively high in this comparison. The plot suggests that hybrid parallel configurations have MPI parallel counterparts with lower degrees of parallelism that consume similar amounts of memory but are faster. The latter are the preferable choice as long as the correspondingly smaller number of nodes offers sufficient memory, at least until further improvements of the hybrid AD performance are in place.

### v2

@@ 4 Benchmarking results and discussion
Throughout Section 3.1, we have changed several aspects of the AD implementation in order to enable hybrid parallel discrete adjoints. To understand the impact of these changes on the overall hybrid AD performance as well as the impact of the optimizations described in Section 3.2, we conduct detailed studies of the single-socket performance on two test cases in Section 4.1. In Section 4.2, we study the impact of the loop parallelization strategies discussed in Section 3.2. Finally, we assess the scalability with a large test case in Section 4.3. In all test cases, we start with converged primal solutions of the RANS equations with the Spalart-Allmaras (SA) turbulence model [42]. The fluid is modelled as an ideal gas. We discretize convective fluxes with the JST scheme [43] and compute spatial gradients with the Green-Gauss approach. See [44] for an overview over the governing equations and numerical schemes. We configure discrete adjoints to compute derivatives of the aerodynamic lift with respect to mesh coordinates, which could, for example, be used for shape optimization. Our goal is to assess comparable computational workloads across different parallel setups, and we therefore always execute a fixed number of discrete adjoint iterations. In particular, we do not make statements about discrete adjoint convergence. Unless mentioned otherwise, we parallelize loops with edge coloring, using the default edge color group size of 512 and adaptivity for discrete adjoints.
@@
Figure 2: Decomposition of the discrete adjoint wall clock time.
We decompose the wall clock time of discrete adjoint computations as visualized in Fig. 2. We distinguish between AD-relevant timings on the one hand, that is, all timings that are directly influenced by the changes described in Sections 3.1 and 3.2, and non-AD timings on the other hand, like preprocessing of the geometry or file I/O. While non-AD timings might be different between MPI-only and hybrid builds of SU2, this is not influenced by hybrid AD itself and not of primary interest in our comparisons. We further refine the AD-relevant timings into recording, evaluation, and management. Recording encompasses the primary and secondary recording passes as explained in Section 2.2 as well as passive computations for the purpose of clearing identifiers. Evaluation refers to all tape evaluations in the course of discrete adjoint iterations (2) or (4), respectively, as well as the final evaluation with respect to parameters in (3). Management encompasses interactions with CoDiPack’s vector of adjoint variables, that is, the seeding prior to tape evaluations as well as the extraction after tape evaluations.
Support for hybrid parallel discrete adjoints is available in the upstream version of SU2. We make a few changes to ensure comparable computational workloads across different parallel setups (e. g., no special behaviour for the case of a single OpenMP thread, lower tolerances so that iterations do not stop early). We use release builds of SU2, compiled with GCC 11.3, with MPI support (OpenMPI 4.0), with AD enabled, with AVX512, and use single precision for linear algebra. The changes made to SU2 and precise build settings are documented in the hybrid_parallel_discrete_adjoints branch [^6] https://github.com/jblueh/SU2/tree/hybrid_parallel_discrete_adjoints (visited on July 16th, 2024), alongside the test cases that are used in the following [^7] https://github.com/jblueh/SU2/tree/hybrid_parallel_discrete_adjoints/benchmark (visited on July 16th, 2024).
@@ 4.1 Single-socket performance
Figure 3: NACA 0012 mesh (left) and Onera M6 mesh (right) for single-socket performance studies.
@@
    We consider the flow around a 2D NACA 0012 airfoil at Mach 0.5 and $0^{\circ}$ angle of attack (AoA), with a Reynolds number (Re) of $20\cdot 10^{6}$. We assume a constant dynamic viscosity of $1.02\cdot 10^{-5}\,\text{Pa\,s}$. The computational domain is discretized by means of a structured mesh with 75140 quadrilaterals, as displayed in the left part of Fig. 3. Given the converged primal flow solution, we perform 300 discrete adjoint iterations in the spirit of (2) followed by one evaluation of (3).
@@
To understand the differences in the AD-specific timings in detail, we consider various intermediate build configurations of SU2 that illustrate the key steps in the transition to hybrid parallel discrete adjoints. We begin with the classical, MPI-only build of SU2 (MPI build, linear). We first exchange the linear identifier management scheme with a reuse scheme (MPI build, reuse). Next, we disable preaccumulations that are incompatible with hybrid parallel AD (MPI build, reuse, hybr. preacc.). We then transition to a hybrid build that uses OpDiLib and features all changes described in Section 3.1 (hybrid build, no opt.). Finally, we enable the optimizations discussed in Section 3.2, first only optimized adjoint vector management (hybrid build, opt. adj. mgmt.) and then in addition also the shared reading optimization (hybrid build, both opt.).
Figure 4: AD-specific performance of the NACA 0012 test case (top) and the Onera M6 test case (bottom). Recording performance (left), management performance (middle), and evaluation performance (right). Serial and parallel timings for various build configurations. Error bars indicate the variation across runs. Speedup factors are relative to the serial run of the respective build.
The left part of Fig. 4 displays the AD recording performance. The respective serial recording timings, for both the NACA 0012 and the Onera M6 test case, illustrate the key differences between the builds. Switching to a reuse management scheme incurs performance overhead due to increased tape size and more involved identifier management, as explained in detail in [34]. Enabling only preaccumulations compatible with hybrid AD reduces the overall number of preaccumulations, which usually reduces the recording time slightly. Switching to a hybrid build requires a thread-safe CoDiPack type, which incurs costs mainly due to tapes accessed via static thread-local pointers [31]. Optimizing for adjoint vector management makes the preaccumulations throughout the recording more efficient, also in the serial case. The shared reading optimization has only a minor impact on the recording performance, the observed small overheads could be due to the tracking of code parts suitable for evaluation without atomics.
@@
The right part of Fig. 4 showcases the AD evaluation performance of the two test cases, which both show similar trends. We explain the evolution of the serial performance. Reuse management leads to larger tapes, both in terms of tape entry size and the overall number of tape entries, and correspondingly larger evaluation times [34]. Likewise, disabling some of the preaccumulations shifts workload from the recording phase to the evaluation phase. The hybrid build’s serial evaluation performance demonstrates the absolute cost of performing atomic updates on adjoint variables. While optimized adjoint vector management does not affect the evaluation performance, the shared reading optimization reduces the overhead by approximately fifty percent in both cases, which underlines the importance of eliminating atomic updates on adjoint variables. We observe consistent MPI speedups of ×$\times 10.3$ to ×$\times 10.9$ and ×$\times 9.3$ to ×$\times 10.4$ for the NACA 0012 and the Onera M6 test case, respectively, and respective OpenMP speedups in the ranges ×$\times 9.5$ to ×$\times 9.9$ and ×$\times 9.8$ to ×$\times 10.0$. As the evaluation inherits the parallelism of the recording, serial parts that limit the OpenMP scaling of the recording will limit the scaling of the evaluation as well. In addition, serialization due to atomic updates on adjoint variables could be a limiting factor for the OpenMP speedup. The observed OpenMP speedups are competitive, which is important for the overall performance as the evaluation time scales with the number of discrete adjoint iterations.
@@
Figure 5: NACA 0012 performance (top) and Onera M6 performance (bottom). Memory high-water marks depending on the type and degree of parallelism for the five SU2 build configurations (left), and joint memory consumption and evaluation time for selected configurations with varying degrees of parallelism (right).
@@
We investigate the memory consumption of the build configurations with varying degrees of parallelism for both test cases in the left part of Fig. 5. We first explain the differences between the build configurations at serial execution. When we compare the memory consumption of hybrid builds to the memory consumption of the classical MPI-only build with linear management, we see that one part of the difference is due to the switch to reuse management, which requires larger individual tape entries and, as copy operations need to be recorded, at the same time more overall tape entries, see [34], where the benefits of optimized copies for SU2 were observed as well. Another part is due to disabling incompatible preaccumulations, which results in larger tapes and therefore higher memory consumption. While both types of overhead have similar magnitudes for the NACA 0012 test case, the additional memory consumption due to disabled preaccumulations becomes dominant in the Onera M6 test case. The optimizations from Section 3.2 do not affect the memory consumption. For each test case, the memory usage increases linearly with an increasing number of MPI processes with similar trends across all build configurations. We attribute this to data duplication across processes, additional halo layers in the partitioning, and larger communication buffers. The memory overhead of increasing the number of OpenMP threads, on the other hand, is significantly smaller, so much that for the NACA 0012 test case, OpenMP parallel execution of hybrid builds requires less per-node memory than MPI parallel execution of the classical build configuration for this test case already at 10-fold parallelism, despite the larger base memory consumption. For the Onera M6 test case, on the other hand, 12-fold parallelism does not suffice to compensate for the memory overhead due to disabled preaccumulations.
@@
We assess the performance tradeoffs in terms of both memory and AD-specific time (aggregated time spent on recording, management, and evaluation), which we plot for selected configurations and their varying degrees of parallelism in the right part of Fig. 5. We consider the classical MPI build with linear identifier management and the fully optimized hybrid build. Clearly, the former dominates the latter in terms of both memory and runtime for the Onera M6 test case. While the hybrid capabilities allow trading memory for runtime at higher degrees of parallelism in the NACA 0012 test case, the MPI build with linear identifier management at 8-fold parallelism still dominates higher degrees of hybrid parallelism. This suggests that, without further improvements, the classical MPI parallelism is the preferred choice for single-socket discrete adjoint computations in SU2. The same recommendation was already made regarding the choice of parallelism for primal single-socket computations in SU2 [30]. In the right part of Fig. 5, we also consider the MPI build with reuse identifier management and only hybrid AD compatible preaccumulations, which is the MPI build that is most similar to the hybrid build in terms of AD and admits a fair assessment of the hybrid AD performance. Starting already at small degrees of parallelism, we see that hybrid AD can serve as a tool to balance memory consumption and runtime.
@@
The results also broaden the perspective on the hybrid parallel discrete adjoint performance. For instance, we expect the notable difference in the memory scaling (left part of Fig. 5) to continue for larger numbers of cores, which could make hybrid parallel discrete adjoints attractive on CPUs with higher core counts. Any principal improvement of hybrid parallel AD such as reduction of the memory offset or improvements of the evaluation time could bridge the small margin between the classical MPI build and the fully optimized hybrid build for the NACA 0012 test case (top right part of Fig. 5). Finally, as the memory overhead in the Onera M6 test case (bottom left part of Fig. 5) is mostly due to disabling preaccumulations, we identify them as a promising target for further improvements of the hybrid parallel discrete adjoint performance.
@@ 4.2 Loop parallelization strategies
We conduct a performance study to support the design decisions regarding loop parallelization for discrete adjoints as outlined in Section 3.2. We study the performance of the Onera M6 test case described in Section 4.1 with both the reduction strategy and edge coloring for loop parallelization, where we vary the edge color group size in the latter. We also assess the performance of the adaptive choice of the edge color group size that we motivated in Section 3.2. The results are displayed in Fig. 6.
Figure 6: Performance of the Onera M6 test case. We compare reductions (violet) to edge coloring with various edge color group sizes (teal and blue). Our adaptive algorithm chooses an edge color group size of 519 (blue). The performance plots show physical memory high-water marks (top left), evaluation time (top right), and recording time (bottom left) together with the variation across multiple runs. The efficiency measure (bottom right) quantifies the imbalance resulting from edge color groups, with an efficiency threshold of 0.875 (red).
@@ 4.3 Scalability
Figure 7: HL-CRM mesh for scaling studies (left) and HL-CRM surface sensitivity (right). The latter is obtained from the derivative of the aerodynamic lift with respect to all mesh coordinates in an additional projection step, treating only surface nodes as design variables.
To assess the scalability of the AD-specific performance, we use an official test case of the Third AIAA CFD High Lift Prediction Workshop involving NASA’s High Lift Common Research Model (HL-CRM) [47], with $8^{\circ}$ AoA at Mach 0.2, $\text{Re}=3.26\cdot 10^{6}$ and a Reynolds length of 275.8 inches. Like in the Onera M6 test case, the dynamic viscosity follows Sutherland’s law with parameters for air. This test case was also used to assess SU2’s hybrid parallel primal performance in [30]. We work with the coarse version of an unstructured mixed-element HL-CRM mesh that consists of approximately 18 million cells and 8.3 million nodes, displayed together with the surface sensitivity in Fig. 7. We start with the converged primal solution and solve Eq. 4 with 1000 GMRES iterations, restarting the Krylov solver every 100th iteration. Afterwards, we evaluate (3) once.
@@
Figure 8: AD-specific performance of the HL-CRM test case. Recording performance (left), management performance (middle), and evaluation performance (right). Parallel timings for various build configurations, ranging from 192-fold to 768-fold parallelism. Error bars indicate the variation across runs. Speedups are relative to the respective timings at 192-fold parallelism.
@@
Fig. 8 showcases the recording, management, and evaluation performance for the HL-CRM test case for the build configurations with various degrees of parallelism. The trends agree with the previous observations in Section 4.1. Considering the recording performance, the relative difference between MPI builds and the hybrid build stands out in this plot. However, this difference is actually similar to or smaller than the corresponding differences in the Onera M6 test case (comparing 12-fold MPI parallelism with 12-fold OpenMP parallelism). Even though the recording is not the major part of the AD-specific runtime, it could be worthwhile to identify and, if possible, improve its parts that do not scale with OpenMP. We see that sufficiently large degrees of parallelism close the gap in the management performance between the hybrid build and the most similar non-hybrid build.
@@
Figure 9: HL-CRM memory high-water marks for the four builds and varying degree of parallelism (left), and joint memory consumption and evaluation time for selected configurations with varying degrees of parallelism (right). Plots display joint memory usage across all involved nodes.
@@
The left part of Fig. 9 displays the memory consumption of the HL-CRM test case with the different build configurations and depending on the degree of parallelism. The plot clearly shows that reuse identifier management does not only introduce an offset in the memory consumption (as observed in the previous test cases) but also affects the scaling of the memory usage. More MPI processes lead to a larger overall number of halo cells and a larger overall number of copy operations for the associated MPI buffers. Unlike recordings with linear identifier management, approaches with reuse identifier management have to record copy operations on the tape, which could explain why their memory usage grows faster with an increasing number of MPI processes. Even though hybrid builds rely on a reuse management scheme as well, introducing 12-fold OpenMP parallelism per MPI process is still sufficient to achieve better memory scaling than the MPI build with linear identifier management. The crossover point lies at approximately 384-fold parallelism for this test case. If we compare the memory consumption of the hybrid configuration with the most similar non-hybrid configuration (MPI, reuse, hybr. preacc.), we see that the hybrid AD approach can save substantial amounts of memory, there is a reduction by a factor of approximately ×$\times 2.4$ at 768-fold parallelism. The plot also shows the additional memory consumption due to disabling preaccumulations, clearly visible as the offset between the curves for MPI builds with reuse management and enabled and disabled preaccumulations, respectively. Re-enabling preaccumulations in the hybrid build could improve its memory consumption up to the same offset. It would shift the crossover point to notably smaller degrees of parallelism and approximately double the memory savings at 768-fold parallelism due to moving from the MPI build with linear management and MPI parallel execution to a hybrid build with hybrid parallel execution in this test case. We expect further improvements on CPUs with larger numbers of cores that admit higher degrees of OpenMP parallelism per MPI process.
@@
The right part of Fig. 9 summarizes the performance in terms of both memory consumption and AD-specific runtime. If we compare the hybrid configuration with the most similar non-hybrid configuration in terms of AD, we see that hybrid AD can be used to trade memory for runtime also in this large test case and correspondingly higher degrees of parallelism. While there are similar tradeoffs between the hybrid configuration and the MPI configuration with linear management, the runtime price for memory reductions is relatively high in this comparison. The plot suggests that hybrid parallel configurations have MPI parallel counterparts with lower degrees of parallelism that consume similar amounts of memory but are faster. The latter are the preferable choice as long as the correspondingly smaller number of nodes offers sufficient memory, at least until further improvements of the hybrid AD performance are in place.